# EEG-Language Modeling for Pathology Detection

### v1

@@ 3.1.1 EEG-language pretraining
$\mathcal{L}_{j,k}^{e2l}$ $=-\log\frac{\exp\left(s^{e2l}_{j,k}/\tau\right)}{\sum_{m=1}^{N}1_{m\neq j}\exp\left(s^{e2l}_{j,m}/\tau\right)}$ (6)
$\mathcal{L}_{j,k}^{l2e}$ $=-\log\frac{\exp\left(s^{l2e}_{j,k}/\tau\right)}{\sum_{m=1}^{N}1_{m\neq j}\exp\left(s^{l2e}_{j,m}/\tau\right)}$ (7)
@@ 3.2.1 Text processing
In order to categorize the textual content in the clinical reports, we employed regular expressions matching for commonly-occurring headings (an overview is provided in Appendix E.2). These enabled the segmentation of individual reports into their respective headings with associated text paragraphs, providing insight into which information in physician reports is encoded in the EEG. We cluster headings into four categories. First, the clinical history cluster contains demographic information in terms of patient age and sex, as well as a brief description of relevant current and/or past pathology and symptoms. The record description cluster includes the physician’s observations of the EEG traces, which describes both normal and abnormal features, often in terms of oscillatory brain activity. The medication cluster contains the patient’s current medication information. Finally, the interpretation cluster summarizes a physician’s thoughts, often including the impression of whether the EEG is normal or pathologically abnormal, as well as a clinical correlation. To investigate whether EEG-language models can learn richer representations by being exposed to a larger variety of text, we also train models by sampling text from all four aforementioned clusters. In order to categorize the textual content in the clinical reports, we employed regular expressions matching for commonly-occurring headings (an overview is provided in Appendix E.2). These enabled the segmentation of individual reports into their respective headings with associated text paragraphs, providing insight into which information in physician reports is encoded in the EEG. We cluster headings into four categories. First, the clinical history cluster contains demographic information in terms of patient age and sex, as well as a brief description of relevant current and/or past pathology and symptoms. The record description cluster includes the physician’s observations of the EEG traces, which describes both normal and abnormal features, often in terms of oscillatory brain activity. The medication cluster contains the patient’s current medication information. Finally, the interpretation cluster summarizes a physician’s thoughts, often including the impression of whether the EEG is normal or pathologically abnormal, as well as a clinical correlation. To investigate whether EEG-language models can learn richer representations by being exposed to a larger variety of text, we also train models by sampling text from all four aforementioned clusters.
@@ 3.2.5 Pathology classification
Notably, while the ELMe,l models trained on either the interpretation cluster or all clusters both perform well with high consistency, training on LLM summaries resulted in highly variable scores. As the LLM-generated text was considerably more uniform with repetitive phrasing across reports, the lack of variability in combination with limited data may have lead to unstable language representations of the text projector. Meanwhile, we observe the opposite pattern for ELMl, where LLM summaries enabled the only consistently high-performing zero-shot classifier. As with the report retrieval analysis, the fixed text representations of a language model which is not finetuned for EEG is likely inadequate to reliably separate between diverse descriptions of pathological and normal EEG. Meanwhile, the rigid LLM-generated text may have aided in this scenario by consistently yielding divergent text representations with which normal and abnormal EEG may be aligned. It must be noted that the current set-up may have favoured the LLM-generated summaries as they were used to subsample the dataset into a balanced subset. However, in this case it is likely this filtering step also aided other models by omitting some EEGs for which the pathology status was uncertain.
@@
ELMl Clinical History 48.42$\pm$1.26 39.54$\pm$1.43 4.92$\pm$1.48
ELMl Medication 50.00$\pm$0.00 76.88$\pm$2.64 0.00$\pm$0.00
ELMl Description 50.00$\pm$0.00 64.57$\pm$5.17 0.00$\pm$0.00
ELMl Interpretation 66.91$\pm$0.95 89.59$\pm$0.27 51.13$\pm$2.27
ELMl LLM Summary 84.32$\pm$0.88 91.00$\pm$0.36 82.45$\pm$1.05
ELMl All Clusters 50.00$\pm$0.00 85.60$\pm$1.03 0.00$\pm$0.00
ELMe,l Clinical History 62.78$\pm$8.58 65.60$\pm$11.31 59.64$\pm$8.15
ELMe,l Medication 53.95$\pm$14.62 55.26$\pm$19.18 52.25$\pm$11.40
ELMe,l Description 64.94$\pm$5.93 68.95$\pm$6.79 62.47$\pm$6.69
ELMe,l Interpretation 82.57$\pm$0.58 91.80$\pm$0.21 80.51$\pm$0.50
ELMe,l LLM Summary 62.42$\pm$15.02 68.13$\pm$22.57 67.24$\pm$8.47
ELMe,l All Clusters 83.61$\pm$0.62 91.79$\pm$0.72 81.62$\pm$0.65
@@ A.1 Optimization
All models are trained using the LARS optimizer [47] with a cosine decay learning rate schedule over 50 epochs, with a warm-up of 4 epochs. The base learning rate is set to 0.3 for EEG-only and 0.01 for EEG-language models, scaled with the batch size (0.2 $\times$ BatchSize/256 [41]). We use a weight-decay parameter of 1 $\times$ 10-4. Models were trained on either an Nvidia Geforce GTX 3090 or Tesla V100 GPU and require less than 24GB of memory. Training took approximately 9 hours for EEG-language modeling or 18 hours for EEG-only modeling due to data augmentations. We used CUDA v12.3 and PyTorch v1.12.1.
@@ A.5 EEG-Only Pretraining
We implement the following three methods for EEG-only SSL:

### v2

@@ 3.1.1 EEG-language pretraining
$\mathcal{L}_{j,k}^{e2l}$ $=-\log\frac{\exp\left(s^{e2l}_{j,k}/\tau\right)}{\sum_{m=1}^{N}\exp\left(s^{e2l}_{j,m}/\tau\right)}$ (6)
$\mathcal{L}_{j,k}^{l2e}$ $=-\log\frac{\exp\left(s^{l2e}_{j,k}/\tau\right)}{\sum_{m=1}^{N}\exp\left(s^{l2e}_{j,m}/\tau\right)}$ (7)
@@ 3.2.1 Text processing
In order to categorize the textual content in the clinical reports, we employed regular expressions matching for commonly-occurring headings (an overview is provided in Appendix E.2). These enabled the segmentation of individual reports into their respective headings with associated text paragraphs, providing insight into which information in physician reports is encoded in the EEG. We cluster headings into four categories. First, the clinical history cluster contains demographic information in terms of patient age and sex, as well as a brief description of relevant current and/or past pathology and symptoms. The record description cluster includes the physician’s observations of the EEG traces, which describes both normal and abnormal features, often in terms of oscillatory brain activity. The medication cluster contains the patient’s current medication information. Finally, the interpretation cluster summarizes a physician’s thoughts, often including the impression of whether the EEG is normal or pathologically abnormal, as well as a clinical correlation. To investigate whether EEG-language models can learn richer representations by being exposed to a larger variety of text, we also train models by sampling text from all four aforementioned clusters.
@@ 3.2.5 Pathology classification
Notably, while the ELMe,l models trained on either the interpretation cluster or all clusters both perform well with high consistency, training on LLM summaries resulted in highly variable scores. As the LLM-generated text was considerably more uniform with repetitive phrasing across reports, the lack of variability in combination with limited data may have lead to unstable language representations of the text projector. Meanwhile, we observe the opposite pattern for ELMl, where LLM summaries enabled the only consistently above-chance zero-shot classifier. As with the report retrieval analysis, the fixed text representations of a language model which is not finetuned for EEG is likely inadequate to reliably separate between diverse descriptions of pathological and normal EEG. Meanwhile, the rigid LLM-generated text may have aided in this scenario by consistently yielding divergent text representations with which normal and abnormal EEG may be aligned. It must be noted that the current set-up may have favoured the LLM-generated summaries as they were used to subsample the dataset into a balanced subset. However, in this case it is likely this filtering step also aided other models by omitting some EEGs for which the pathology status was uncertain.
@@
ELMl Clinical History 50.00$\pm$0.00 34.34$\pm$1.66 0.00$\pm$0.00
ELMl Medication 50.00$\pm$0.00 74.99$\pm$1.80 0.00$\pm$0.00
ELMl Description 50.00$\pm$0.00 43.70$\pm$1.99 0.00$\pm$0.00
ELMl Interpretation 50.00$\pm$0.00 89.23$\pm$0.31 0.00$\pm$0.00
ELMl LLM Summary 71.98$\pm$0.62 90.92$\pm$0.35 61.77$\pm$1.24
ELMl All Clusters 50.00$\pm$0.00 84.52$\pm$0.57 0.00$\pm$0.00
ELMe,l Clinical History 62.50$\pm$8.74 67.57$\pm$10.41 60.96$\pm$6.13
ELMe,l Medication 51.09$\pm$17.79 52.21$\pm$23.54 50.67$\pm$12.53
ELMe,l Description 64.03$\pm$3.95 71.40$\pm$4.32 63.56$\pm$3.81
ELMe,l Interpretation 82.34$\pm$1.42 91.80$\pm$0.47 80.10$\pm$1.63
ELMe,l LLM Summary 58.87$\pm$15.48 67.98$\pm$20.88 64.32$\pm$9.06
ELMe,l All Clusters 83.16$\pm$1.15 91.91$\pm$0.67 81.25$\pm$1.35
@@ A.1 Optimization
All models are trained using the LARS optimizer [47] with a cosine decay learning rate schedule over 50 epochs, with a warm-up of 4 epochs. The base learning rate is set to 0.3 for EEG-only and 0.01 for EEG-language models, scaled with the batch size (BaseLR $\times$ BatchSize/256 [41]). We use a weight-decay parameter of 1 $\times$ 10-4. Models were trained on either an Nvidia Geforce GTX 3090 or Tesla V100 GPU and require less than 24GB of memory. Training took approximately 9 hours for EEG-language modeling or 18 hours for EEG-only modeling due to data augmentations. We used CUDA v12.3 and PyTorch v1.12.1.
@@ A.5 EEG-Only Pretraining
We implement the following methods for EEG-only SSL: